OpenStudy (anonymous):

(x 3 - 4x^ 2 + 4x - 1) ÷ (x -1)

OpenStudy (anonymous):

Do you know how to do synthetic division?

OpenStudy (anonymous):

no :c

OpenStudy (anonymous):

Here is a 5 minute video to help you then! https://www.khanacademy.org/math/algebra2/polynomial_and_rational/synthetic-division/v/synthetic-division

OpenStudy (anonymous):

Do they want you to do long division or synthetic though? You'll need to know both ways.

OpenStudy (anonymous):

Focus on one for now.

OpenStudy (anonymous):

Tell me when you start, so I can check your work.

OpenStudy (campbell_st):

here is a simple method that uses grouping in pairs \[\frac{(x^3 -1) - (4x^2 - 4x)}{x -1}\] now factor the numerator (x^3 -1) is the difference of 2 cubes and can be factored to \[x^3 -1 = (x -1)(x^2 -x + 1)\] and \[4x^2 - 4x = 4x(x -1)\] so the problem now becomes \[\frac{x^3 - 4x^2 + 4x -1}{x -1} = \frac{(x -1)(x^2 - x + 1) -4x(x - 1)}{(x -1)}\] now remove the common factror and simplify for the answer hope it helps